OpenStudy (anonymous):

Determine the solution of the system. y = x + 6 y = 2x

OpenStudy (anonymous):

2x = x +6

OpenStudy (anonymous):

Plug in 2x for y since that is the value. y = x + 6 2x = x + 6 x = 6 Plug this in for x y = 2x y = 2(6) y = 12 (6, 12

OpenStudy (anonymous):

Solve for x, then plug that value into either equation to obtain y value

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

x=6, y=12